Find two arithmetic means between -17 and -71.
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

= (-17 + (-71))/2
= -88/2
= -44

Answer by Edwin McCravy(20054)


The other tutor's solution is incorrect.  He only inserted 
one arithmetic mean.

Find two arithmetic means between -17 and -71.

-17, ___, ___, -71

So the first term, a1 is -17 and the 4th term a4 is -71.

an = a1 + (n - 1)d

We substitute n=4, 

a4 = a1 + (4 - 1)d
a4 = a1 + 3d

We substitute a1=-17 and a4=-71

-71 = -17 + 3d

Solve for d

-54 = 3d
-18 = d

-17, ___, ___, -71

To fill in the first blank -17+(-18) = -35  

-17, -35, ___, -71

To fill in the second blank -35+(-18) = -53

-17, -35, -53, -71

That's it.  But as a check let's add -18 to -53.
Yes, we get (-53)+(-18) = -71.

So the two arithmetic means are -35 and -53.

Edwin